4. Find the third, sixth, and eighth terms of the sequence A(n) = 5 + (n – 1)(3).
Answer by drk(1908)

We have the following question:
Find the third, sixth, and eighth terms of the sequence A(n) = 5 + (n – 1)(3).
To find the terms, let n = 3, 6, and 8 using
A(3) = 5 + (3 – 1)*3 = 5 + 2*3 = 11
A(6) = 5 + (6 – 1)*3 = 5 + 5*3 = 20
A(8) = 5 + (8 – 1)*3 = 5 + 8*3 = 26